Igi is now twice as old as Bill. Seven years ago, the sum of their ages was 16. How old are they now?
Answer by monika_p(71)

x- Igi's age
y-Bill's age
x=2y info from the first sentence
(x-7)+(y-7)=16 from the second sentence
substitute for x
(2y-7)+(y-7)=16
3y-14=16
3y=30
y=10 then x=20